OpenStudy (anonymous):

x=vt cosθ and y=- 1/2 gt^2+vt sinθ+h find the Cartesian equation of motion

OpenStudy (perl):

solve for t

OpenStudy (perl):

x/( vcos theta) = t y = -1/2 g [ x/( vcos theta) ]^2 + v[x/( vcos theta) ] sin theta + h

OpenStudy (perl):

y = -1/2 * g * x^2 / ( v^2 cos^2 theta) + x* tan(theta) + h

OpenStudy (perl):

y = -g/[2v^2 cos^2 theta]*x^2 + tan(theta)* x + h

OpenStudy (anonymous):

OK thankyou very much perl